How do I write a quadratic equation with the given roots (2/3) and -4 in the form of ax^2+bx+c=0
Answer by ankor@dixie-net.com(22740)

write a quadratic equation with the given roots (2/3) and -4 in the form of ax^2+bx+c=0
:
Derive the factors from the given roots
:
x = 2/3
(3x - 2) = 0
and
x = -4
(x + 4) = 0
:
FOIL: (3x-2)(x+4)
3x^2 + 12x - 2x - 8
y = 3x^2 + 10x - 8, is the equation